OpenStudy (anonymous):

revolve the region in first quadrant bounded by y = x^2 , y = x , about line x = 2, using washer method

OpenStudy (anonymous):

\[\pi \int\limits_{0}^{4}(2-x)^{2}-(2-x ^{2})^{2}\]

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

noooooooooo

OpenStudy (anonymous):

srry its been a long time.

OpenStudy (anonymous):

the region is from (0,0) to (1,1)

OpenStudy (anonymous):

oh sorry sorry forgot about the bounds. assumed that x22 was also a bound

OpenStudy (anonymous):

also youre using dy , not x

OpenStudy (anonymous):

ok nvm i rambl.ed

OpenStudy (anonymous):

now you will die

OpenStudy (anonymous):

ure supposed to use dx?

OpenStudy (anonymous):

dy

OpenStudy (anonymous):

i was using DY

OpenStudy (anonymous):

since its horizontal integration

OpenStudy (anonymous):

just the wrong values

OpenStudy (anonymous):

no its wrong in other ways!!!

OpenStudy (anonymous):

ok lets do it again: \[\pi \int\limits_{0}^{1} (2-x)^{2}-(2-x ^{2})^{2}\]

OpenStudy (anonymous):

this HASSSSS TO BE RIGHT

OpenStudy (anonymous):

eh eh?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its pi* integral ( y - 2)^2 - (sqrt y - 2)^2